the base of a
isosceles triangle is 12 cm and
its perimeter is 32 cm.find the
area of the triangle?

Answer by lwsshak3(11628)

the base of a
isosceles triangle is 12 cm and
its perimeter is 32 cm.find the
area of the triangle?
***
Area of a triangle=1/2*base*height
The two equal sides of the isoceles triangle=(32-12)/2=20/2=10
height=vertical line from apex to the base=√(10^2-6^2)=√(100-36)=√64=8
Area=1/2*12*8=48 cm^2